Cube root of 64a^6 b^2
Answer by stanbon(75887)

Cube root of 64a^6 b^2
-----
Cube root of 2^6 * a^6* b^2 = 2^2*a^2*b^(2/3) = 4a^2b^(2/3)
---------
Cheers,
Stan H.
----------